High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates:

Department
Molding Painting
Direct labor-hours 34,000 51,400
Machine-hours 90,000 40,000
Fixed manufacturing overhead cost $ 171,000 $ 457,460
Variable manufacturing overhead per machine-hour $ 3.00 -
Variable manufacturing overhead per direct labor-hour - $ 5.00

Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job:

Department
Molding Painting
Direct labor-hours 87 133
Machine-hours 360 74
Direct materials $ 948 $ 1,160
Direct labor cost $ 740 $ 970

Required:

1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department.

2. Compute the total overhead cost applied to Job 205.

3-a. What would be the total manufacturing cost recorded for Job 205?

3-b. If the job contained 33 units, what would be the unit product cost?
Answer #1
1
Predetermined overhead rate:
Molding Department 4.90 per MH =3+(171000/90000)
Painting Department 13.90 per DLH =5+(457460/51400)
2
Overhead applied:
Molding Department 1764 =360*4.90
Painting Department 1849 =133*13.90
Total overhead cost 3613
3a
Molding Painting Total
Direct materials 948 1160 2108
Direct labor 740 970 1710
Manufacturing Overhead 1764 1849 3613
Total cost 3452 3979 7431
3b
Total cost 7431
Divide by units 33
Unit product cost 225.18 per unit